2. Genius Inc., obtained significant influence over K Corp by acquiring 30% of K Corp's 100,000 outstanding ordinary shares at a total cost of $5 per share on January 1, 2018. On June 15, K Corp declared and paid a cash dividend of $100,000. On December 31, K Corp. reported a net income of $250,000 for the year.

Instructions

Prepare all the necessary journal entries for 2018 for Genius Inc and compute the balance in the investment account in K Corp on December 31, 2018
Answer #1
Debit Credit
January 1, 2018 Investment in shares of K Corp 500000 =100000*5
       Cash 500000
June 15 Cash 30000 =100000*30%
     Investment in shares of K Corp 30000
December 31 Investment in shares of K Corp 75000 =250000*30%
      Revenue from Investment 75000
Balance in investment account in K Corp 545000 =500000+75000-30000